# Closing the Data-Efficiency Gap Between Autoregressive and Masked Diffusion LLMs

### v2

@@ A.2 LLM usage
The usage of LLM is limited to language polishing and literature search. We asked an LLM to suggest surface-level rewrites to improve clarity, grammar, and style for author-written passages. Edits were limited to phrasing and organization at the sentence/paragraph level. We also used an LLM to source papers, and produce brief literature summaries for writing references.
@@
We construct a Wiki from real Wikipedia articles following the protocol of pan2025memorization. We first crawl all the pages under the wiki category “Category:2025_by_month”, then filter out the pages that are created before January 1st, 2025. This process minimizes the leakage of this “new” knowledge to the base model. Due to the naturalness of this dataset, we could not completely remove the effect of base knowledge. Llada-Instruct has a slightly higher base model accuracy than Llama-3.1-8B-instruct, but they are qualitatively similar (Table 1). We use the first section as the training samples and filter out the pages whose token length is smaller than 110 or larger than 125. This results in 96 wiki articles. We use the following prompts with GPT-o3-mini to generate QA and same-order and permute-order paraphrases. We classify QAs into forward and backward styles. This is done by prompting GPT-o3-mini to generate keywords in the question and answer, then comparing their appearance order in the original text.
@@
Your task is to paraphrase a text paragraph. The paragraph is given below. Make sure to keep the same meaning but change the wording. Do not change any factual information. Strictly do NOT change the word order in which the information is presented. Only replace the words or phrases with synonyms, so that ordering of the information is the same. Try to keep roughly the same length of the original text. Give 9 different paraphrases for each text. Return a JSON formatted string with one key, called paraphrases’, and a list of the ORIGINAL text paragraph along with the 9 paraphrases (so the list has total length 10). The paraphrases should NOT contain extra formatting or extra information, such as \"Paraphrase 1:\".
{passage}
@@
Your task is to paraphrase a text paragraph. The paragraph is given below. Make sure to keep the same meaning but change the wording. Do not change any factual information. Change the word order in which the information is presented. Think about the order in three levels: word, sentence, and paragraph.
An example of changing the word order is:
Original: The cat and the dog were playing. Paraphrase: The dog and the cat were playing.
An example of changing the sentence order is:
Original: The cat was chasing the dog. Paraphrase: The dog was being chased by the cat.
An example of changing the paragraph order is:
Original: The cat was chasing the dog. Then, the cat got tired. Paraphrase: The cat got tired. Before that, the cat was chasing the dog.
Try to keep roughly the same length of the original text. Give 9 different paraphrases for each text. Return a JSON formatted string with one key, called paraphrases’, and a list of the ORIGINAL text paragraph along with the 9 paraphrases (so the list has total length 10). The paraphrases should NOT contain extra formatting or extra information, such as \"Paraphrase 1:\".
{passage}
@@
Your task is to generate several question, answer, and cue used in the question triplets based on a given passage below. Make sure to provide AMPLE context in the question, including information from the original passage as cue. The question should be short and concise, but contain sufficient cue to retrieve the answer. Do not use pronouns in the question. Use the exact words from the passage as the cue. The questions will be used for a close-book test. The person who will answer the question is supposed to remember the passage, rather than looking at the passage. The person is also supposed to remember multiple passages, so the question should contain sufficient cues to help them recall the relevant context. Do not mention according to the passage’, or other redundant wordings. Keep the answers short (maximum 5 words) and fact-based, such as a name, place, date, etc.. Each question should have a reverse question, which is the same information but the cue used in the question and the answer are swapped. For example, if the question is What is the capital of France?’, the reverse question should be Paris is the capital of which country?’.
Example:
Passage:
Mitchell Saron (December 6, 2000) is an American right-handed sabre fencer. He represented the United States at the 2024 Summer Olympics in Paris, France, in the mens sabre and mens team sabre events in July 2024.
Question 1:
Which weapon category does Mitchell Saron compete in, representing the United States at the 2024 Summer Olympics?
Answer 1:
Sabre
Cue used in the question:
[Mitchell Saron, United States, 2024 Summer Olympics]
Question 2 (reverse question of question 1):
Who represented the United States at the 2024 Summer Olympics to compete in the mens sabre?
Answer 2:
Mitchell Saron
Cue used in the question:
[Sabre, United States, 2024 Summer Olympics]
Return a JSON formatted string with one key, called qa_data’, and a list of (question, answer, cue_used_in_question) tuples. Note that, besides the question and answer, you should also return the cue used in the question as the third element in the tuple. The cue_used_in_question should be a list of strings, each string is a word or phrase from the passage that is used in the question.
Passage:
{passage}
@@
All the training and inference code will be available in an online repository. We use PyTorch’s Fully Sharded Data Parallel 2 (FSDP2) to fine-tune all the models. We find that using mixed precision training is important for the fine-tuning performance (around 30% performance gain), and use the configs: MixedPrecisionPolicy(param_dtype="bf16", reduce_dtype="float32", cast_forward_inputs=True). All the experiments are full parameter fine-tuning on 4x 80G H100 GPUs. We use a batch size of 64 (16 per device) for all the experiments. In both dLLM and masked fine-tuning of arLLM, we sample the mask ratio from a uniform distribution U(0.05,0.95) for each batch (except for the fixed mask ratio experiments). Note that, unlike the original dLLM training recipes which use U(0,1) (llada), given that our sequence length is much shorter than the pre-training, we leave a small margin to avoid edge cases.
While doing masked fine-tuning of Llama models, we pick a reserved special token whose token id is 128013 in the Llama tokenizer. While doing masked fine-tuning of Qwen models, we pick token “[]” whose token id is 1294 in the Qwen tokenizer.
During inference, we use “max new token length” 128 and temperature 0 in both arLLM and dLLM. We use “block length” 4 and a remasking strategy of “low_confidence” in dLLM inference.
We use Adam optimizer with 0.1 weight decay coefficient; betas 0.9 and 0.95; 2% total steps as warm-up steps. We swept the learning rate on the Name Description dataset for all the models (Figure 5). We choose to use learning rates that yield smooth gains in accuracy throughout the training process while achieving high final accuracy. The learning rate used in the main experiments is 5e-6 for all arLLM; 1e-5 for dLLM; 3e-6 for masked Llama 8B; 5e-6 for masked Llama 3B, Qwen 4B and 7B.
@@
A.5 Masked SFT configs
In Section 8, we proposed a Masked SFT methods that turns any SFT prompt-response sequence into a demasking task. Specifically, the constructed demasking task sequence contains the question and and masked answer in the user prompt, and full answer in the assistant response. For example, an data entry in th GSM8K dataset is: Question: Joy can read 8 pages of a book in 20 minutes. How many hours will it take her to read 120 pages? Answer: In one hour, there are 3 sets of 20 minutes. So, Joy can read 8 x 3 = <<8*3=24>>24 pages in an hour. It will take her 120/24 = <<120/24=5>>5 hours to read 120 pages. 5. After generating a random mask, the constructed masked SFT sequence is:
[…]
The highlighted tokens are where the loss is calculated. More specifically, in Eq. 2, $m_{i}$ is set to $1$ when the $i$th token is in the assistant response and its corresponding token in the user prompt is masked.
We choose to use the GSM8K and MATH datasets for testing the SFT performance on Llama-3.2-3B-Instruct and Qwen/Qwen3-4B-Instruct-2507. For the MATH dataset, we use the default subset from huggingface DigitalLearningGmbH/MATH-lighteval. We further filter out the training samples whose token lengths (question + answer) are longer than 512. When evaluating the resulting models, we use the evaluation framework LM Evaluation Harness and the default tasks gsm8k and hendrycks_math (eval-harness). Specifically, we choose to use 0-shot and pass@1 with a maximum generation length of 256 at a temperature of 0. For GSM8K we report the accuracy using exact match with LM Evaluation Harness’s flexible extraction. For MATH we report the accuracy using exact match with math-verify extraction (Kydlicek_Math-Verify_Math_Verification). Both extraction methods are chosen to maximize alignment with human examination.
@@ A.5 Masked SFT configs
A.6 On reversal curse
@@
Prior studies have justified the reversal curse as an intrinsic limitation of arLLM training (zhu2024towards; kitouni2024factorization; zhu2024towards). Here we provide an explanation that is conceptually easy to grasp. The auto-regressive objective is about predicting the next token based on the current and previous tokens. If the prediction of one next token requires a piece of new knowledge (i.e., it cannot be predicted based on the current knowledge in the weights or previous tokens), the loss will force the weights to change to favor such a prediction. More specifically, the change of weights induces a different representation (i.e., intermediate layer activations) of the previous tokens that favors the prediction of the next token. Since feedforward layers can be considered associative memory (meng2022locating), the change, conceptually, could be associating a new attribute with the representation of a token. However, such change does not affect the representation of future tokens to favor the prediction of the current token, since they do not contribute to the prediction of the “next” token. Thus, the future tokens could not learn a new association to its preceeding tokens. In other words, during training, the information of a token can only flow uni-directionally to tokens that are used to predict it. This has been named the “factorization curse" (kitouni2024factorization). This intuition can explain why the masked fine-tuning of arLLM resolves the curse. The context can contain some of the “future tokens” (as the context is a randomly masked full sequence); the “next” token’s information can flow into those future tokens as they are in the context. Paraphrases mitigates the reversal curse by creating views where the “future” cue becomes left context. For a fact $(x\!\to\!y)$ usually realized as “xy…x…y…”, paraphrases are fluent variants that put y before x (passive voice, topicalization, apposition, clause reordering), producing complementary left-to-right gradients and corpus-level bidirectional coverage without changing the AR objective. However, globally reversing all the information units in the documents can breaks discourse, coreference, and grammar, creating distribution shift.
A.7 Generation eamples

### v3

@@ A.2 LLM usage
The usage of LLM is limited to language polishing and literature search. We asked an LLM to suggest surface-level rewrites to improve clarity, grammar, and style for author-written passages. Edits were limited to phrasing and organization at the sentence/paragraph level. We also used an LLM to source papers and produce brief literature summaries for writing references.
@@ A.3 Dataset details and examples
We construct a Wiki dataset from real Wikipedia articles following the protocol of (Pan et al., 2025). We first crawl all the pages under the wiki category “Category:2025_by_month”, then filter out the pages that are created before January 1st, 2025. This process minimizes the leakage of this “new” knowledge to the base model. Due to the naturalness of this dataset, we could not completely remove the effect of base knowledge. Llada-Instruct has a slightly higher base model accuracy than Llama-3.1-8B-instruct, but they are qualitatively similar (Table 1). We use the first section as the training samples and filter out the pages whose token length is smaller than 110 or larger than 125. This results in 96 wiki articles. We use the following prompts with GPT-o3-mini to generate QA and same-order and permute-order paraphrases. We classify QAs into forward and backward styles. This is done by prompting GPT-o3-mini to generate keywords in the question and answer, then comparing their appearance order in the original text.
@@
Prompt for generating same-order paraphrases Your task is to paraphrase a text paragraph. The paragraph is given below. Make sure to keep the same meaning but change the wording. Do not change any factual information. Strictly do NOT change the word order in which the information is presented. Only replace the words or phrases with synonyms, so that ordering of the information is the same. Try to keep roughly the same length of the original text. Give 9 different paraphrases for each text. Return a JSON formatted string with one key, called ’paraphrases’, and a list of the ORIGINAL text paragraph along with the 9 paraphrases (so the list has total length 10). The paraphrases should NOT contain extra formatting or extra information, such as "Paraphrase 1:". {passage}
@@
Prompt for generating permute-order paraphrases Your task is to paraphrase a text paragraph. The paragraph is given below. Make sure to keep the same meaning but change the wording. Do not change any factual information. Change the word order in which the information is presented. Think about the order in three levels: word, sentence, and paragraph. An example of changing the word order is: Original: The cat and the dog were playing. Paraphrase: The dog and the cat were playing. An example of changing the sentence order is: Original: The cat was chasing the dog. Paraphrase: The dog was being chased by the cat. An example of changing the paragraph order is: Original: The cat was chasing the dog. Then, the cat got tired. Paraphrase: The cat got tired. Before that, the cat was chasing the dog. Try to keep roughly the same length of the original text. Give 9 different paraphrases for each text. Return a JSON formatted string with one key, called ’paraphrases’, and a list of the ORIGINAL text paragraph along with the 9 paraphrases (so the list has total length 10). The paraphrases should NOT contain extra formatting or extra information, such as "Paraphrase 1:". {passage}
@@
Prompt for generating QAs Your task is to generate several question, answer, and cue used in the question triplets based on a given passage below. Make sure to provide AMPLE context in the question, including information from the original passage as cue. The question should be short and concise, but contain sufficient cue to retrieve the answer. Do not use pronouns in the question. Use the exact words from the passage as the cue. The questions will be used for a close-book test. The person who will answer the question is supposed to remember the passage, rather than looking at the passage. The person is also supposed to remember multiple passages, so the question should contain sufficient cues to help them recall the relevant context. Do not mention ’according to the passage’, or other redundant wordings. Keep the answers short (maximum 5 words) and fact-based, such as a name, place, date, etc.. Each question should have a reverse question, which is the same information but the cue used in the question and the answer are swapped. For example, if the question is ’What is the capital of France?’, the reverse question should be ’Paris is the capital of which country?’. Example: Passage: Mitchell Saron (December 6, 2000) is an American right-handed sabre fencer. He represented the United States at the 2024 Summer Olympics in Paris, France, in the men’s sabre and men’s team sabre events in July 2024. Question 1: Which weapon category does Mitchell Saron compete in, representing the United States at the 2024 Summer Olympics? Answer 1: Sabre Cue used in the question: [Mitchell Saron, United States, 2024 Summer Olympics] Question 2 (reverse question of question 1): Who represented the United States at the 2024 Summer Olympics to compete in the men’s sabre? Answer 2: Mitchell Saron Cue used in the question: [Sabre, United States, 2024 Summer Olympics] Return a JSON formatted string with one key, called qa_data, and a list of (question, answer, cue_used_in_question) tuples. Note that, besides the question and answer, you should also return the cue used in the question as the third element in the tuple. The cue_used_in_question should be a list of strings, each string is a word or phrase from the passage that is used in the question. Passage: {passage}
@@ A.4 Training configs
All the training and inference code will be available in an online repository. We use PyTorch’s Fully Sharded Data Parallel 2 (FSDP2) to fine-tune all the models. We find that using mixed precision training is important for the fine-tuning performance (around 30% performance gain), and use the configs: MixedPrecisionPolicy(param_dtype="bf16", reduce_dtype="float32", cast_forward_inputs=True). All the experiments are full parameter fine-tuning on 4x 80G H100 GPUs. We use a batch size of 64 (16 per device) for all the experiments. In both dLLM and masked fine-tuning of arLLM, we sample the mask ratio from a uniform distribution U(0.05,0.95) for each batch (except for the fixed mask ratio experiments). Note that, unlike the original dLLM training recipes which use U(0,1) (Nie et al., 2025b), given that our sequence length is much shorter than the pre-training, we leave a small margin to avoid edge cases.
While doing masked fine-tuning of Llama models, we use a reserved special token whose token id is 128013 in the Llama tokenizer. While doing masked fine-tuning of Qwen models, we pick token “[]” whose token id is 1294 in the Qwen tokenizer.
During inference, we use “max new tokens” 128 and temperature 0 in both arLLM and dLLM. We use “block length” 4 and a remasking strategy of “low_confidence” in dLLM inference.
We use Adam optimizer with 0.1 weight decay coefficient; betas 0.9 and 0.95; 2% total steps as warm-up steps. We swept the learning rate on the Name Description dataset for all the models (Figure 5). We choose learning rates that yield smooth accuracy gains and high final accuracy. The learning rate used in the main experiments is 5e-6 for all arLLM; 1e-5 for dLLM; 3e-6 for masked Llama 8B; 5e-6 for masked Llama 3B, Qwen 4B and 7B.
@@
A.5 Masked SFT training configs
In Section 4.5, we proposed a Masked SFT method that turns any SFT prompt-response sequence into a demasking task. Specifically, the constructed demasking task sequence contains the question and masked answer in the user prompt, and full answer in the assistant response. For example, a data entry in the GSM8K dataset is: Question: Joy can read 8 pages of a book in 20 minutes. How many hours will it take her to read 120 pages? Answer: In one hour, there are 3 sets of 20 minutes. So, Joy can read 8 x 3 = <<8*3=24>>24 pages in an hour. It will take her 120/24 = <<120/24=5>>5 hours to read 120 pages. 5. After generating a random mask, the constructed masked SFT sequence is:
We compute loss only on the highlighted tokens. More specifically, in Eq. 2, $m_{i}$ is set to $1$ when the $i$th token is in the assistant response and its corresponding token in the user prompt is masked.
We choose to use the GSM8K and MATH datasets for testing the SFT performance on Llama-3.2-3B-Instruct and Qwen/Qwen3-4B-Instruct-2507. For the MATH dataset, we use the default subset from huggingface DigitalLearningGmbH/MATH-lighteval. We further filter out the training samples whose token lengths (question + answer) are longer than 512. When evaluating the resulting models, we use the evaluation framework LM Evaluation Harness and the default tasks gsm8k and hendrycks_math ((Gao et al., 2024)). Specifically, we choose to use 0-shot and pass@1 with a maximum generation length of 256 at a temperature of 0. For GSM8K we report the accuracy using exact match with LM Evaluation Harness’s flexible extraction. For MATH we report the accuracy using exact match with math-verify extraction ((Kydlíček, )). Both extraction methods are chosen to maximize alignment with human examination.
@@ A.5 Masked SFT training configs
A.6 Compute overhead analysis
We use the Wiki dataset to comprehensively characterize the computational cost of different training methods tested in Table 1 covering data preparation, training, inference and convergence to peak accuracy. Specifically, we test 8b parameter arLLM (Llama-3.1-8b-instruct) and dLLM (Llada) models in the following 5 conditions: (1) arLLM + pretraining style fine-tuning + w/o paraphrases, (2) arLLM + pretraining style fine-tuning + w paraphrases, (3) arLLM + pretraining style fine-tuning + reverse training (Golovneva et al., 2024b), (4) dLLM + pretraining style fine-tuning + w/o paraphrases, (5) arLLM + masked SFT style fine-tuning + w/o paraphrases (ours) (Table 4).
Data Preparation
Only condition (2) requires computationally expensive data preparation to generate semantically identical and natural paraphrases of the original dataset. For example, in the Wiki dataset, paraphrasing 10 per sample costs 0.1M generation tokens of GPT-o3-mini or equivalent models. This cost scales further with larger original datasets and the requirements for the diversity of paraphrases. In contrast, all other methods require no or very simple data transformations during the training (sampling mask or reverse sequence) without additional compute.
Training
We compare training FLOPs, wall time and peak allocated memory per step. Theoretical FLOPs calculation follows formulation introduced in (Kaplan et al., 2020). Specifically, the training cost of a dense transformer is approximated as $2N$ FLOPs per token for the forward pass and $4N$ for backpropagation, in total $6N$ FLOPs per token, where $N$ denotes the model parameter size. For a step that processes $T$ tokens in total, the theoretical per-step cost is therefore $\mathrm{FLOPs/step}\;\approx\;6\,N\,T$. In our comparisons, we keep the model size fixed at N=8N{=}8B and use the same effective global batch size across all conditions, so the only varying factor is the average sequence length of training samples. We thus report the theoretical quantity as a factor of $T$. Since paraphrasing or reverse training only modify the information order within the training samples, condition (1)-(4) have comparable average sequence length $T$. Condition (5) requires presenting the model with the original sequence and its masked counterpart in SFT format (Figure 3) doubling the average sequence length to $2T+c$ where $c$ is the constant accounting for SFT instructions. We also empirically measure the average FLOPs, wall time and peak memory per step using PyTorch Profiler. Empirical FLOPs measurements agree with the theoretical calculations. Among all conditions, our proposed method condition (5) costs around twice training FLOPs and slightly larger wall time as well as peak allocated memory. Additionally, we report that the sampling the mask and constructing the masked fine-tuning prompt during training for condition (4) and (5) takes negligible time (<3.5ms) and memory compared to the forward/backward pass, thus causing no significant overhead.
Inference
For inference statistics, we only compare the theoretical inference FLOPs as empirical values highly depend on the generation length of a particular answer. Following the same calculation detailed in the above section, an inference step produces a forward pass through the dense transformer, leading to $\mathrm{FLOPs/step}\;\approx\;2\,N\,S$ where $S$ denotes the average generation sequence length. Only condition (4) dLLM needs inference FLOPs quadratic in $S$ as generation in dLLM cannot reuse the KV cache since it changes after each denoising step (Ni et al., 2025). All other conditions are linear in $S$.
Convergence
[…]
A.7 On reversal curse
@@
Prior studies have justified the reversal curse as an intrinsic limitation of arLLM training (Zhu et al., 2024; Kitouni et al., 2024; Zhu et al., 2024). Here we provide an explanation that is conceptually easy to grasp. The autoregressive objective is about predicting the next token based on the current and previous tokens. If the prediction of one next token requires a piece of new knowledge (i.e., it cannot be predicted based on the current knowledge in the weights or previous tokens), the loss will force the weights to change to favor such a prediction. More specifically, the change of weights induces a different representation (i.e., intermediate layer activations) of the previous tokens that favors the prediction of the next token. Since feedforward layers can be considered associative memory (Meng et al., 2022), the change, conceptually, could be associating a new attribute with the representation of a token. However, such change does not affect the representation of future tokens to favor the prediction of the current token, since they do not contribute to the prediction of the “next” token. Thus, the future tokens could not learn a new association to its preceding tokens. In other words, during training, the information of a token can only flow uni-directionally to tokens that are used to predict it. This has been named the “factorization curse” (Kitouni et al., 2024). This intuition can explain why the masked fine-tuning of arLLM resolves the curse. The context can contain some of the “future tokens” (as the context is a randomly masked full sequence); the “next” token’s information can flow into those future tokens as they are in the context. Paraphrases mitigate the reversal curse by creating views where the “future” cue becomes left context. For a fact $(x\!\to\!y)$ usually realized as “xy…x…y…”, paraphrases are fluent variants that put y before x (passive voice, topicalization, apposition, clause reordering), producing complementary left-to-right gradients and corpus-level bidirectional coverage without changing the AR objective. However, globally reversing all the information units in the documents can break discourse, coreference, and grammar, creating distribution shift.
A.8 Generation examples